2. Adam is the owner of a small grocery store in a busy section of Boulder, Colorado. Adam’s annual revenue is $200,000 and his total explicit cost (Adam pays himself an annual salary of $30,000) is $180,000 per year. A supermarket chain wants to hire Adam as its general manager for $60,000 per year.

a. What is the opportunity cost to Adam of owning and managing the grocery store?
b. What is Adam’s accounting profit?

c. What is Adam’s economic profit
1
Expert's answer
2020-10-25T19:27:55-0400

(a) Opportunity cost is the $60,000 in foregone salary that Adam might have earned had he decided to work as a general manager for the supermarket chain.

(b) TR = Total revenue

Tc =total cost

"\pi=" profit

Adam's accounting profit

"\pi=TR - TC_{explicit}=" "\$200, 000 - \$180, 000 = \$20, 000"

(c) Adam's economic profit

"\pi = TR - TC_{explicit} - TC _{implicit}= \$200, 000 - \$180, 000 - \$30, 000= -\$10000"